OpenStudy (anonymous):

Integration Just to double check my answer @ganeshie8 @hartnn

OpenStudy (anonymous):

Given\[y=\frac{ 1 }{ 2x-3 }\]find dy/dx.Hence,find\[\int\limits_{}^{}\frac{ 1 }{ (2x-3)^2 }dx\]

hartnn (hartnn):

ok, whats your dy/dx?

OpenStudy (anonymous):

\[\frac{ dy }{ dx }=\frac{ -2 }{ (2x-3)^2 }\]

hartnn (hartnn):

thats correct. and the answer to the final integration that you got ?

OpenStudy (anonymous):

i got\[\int\limits_{}^{}\frac{ 1 }{ (2x-3)^2 }dx=\frac{ 1 }{ 6-4x }\]

hartnn (hartnn):

i get the same :) \(\checkmark \)

OpenStudy (anonymous):

Thank you @hartnn :)

hartnn (hartnn):

welcome ^_^